OpenStudy (anonymous):

A mass weighing 2 lbs stretches a spring 6 inches. At t=0 the mass is released from a point 8 inches below the equilibrium position with an upward velocity of 4/5ft/s. Determine the equation of free motion.

OpenStudy (anonymous):

Stuck. I don't get how to write the equation of free motion

OpenStudy (anonymous):

not sure if it means: Acos(wt-d) or m*d^2/dt^2 +k/m*y=0

OpenStudy (irishboy123):

the latter, the differential equation, is the equation of free motion. the former is a standard solution

OpenStudy (anonymous):

ah so I just need to plug in the values? Not sure where to plug in though ... :/

OpenStudy (anonymous):

is x'(0) = 4/5ft/s?

OpenStudy (irishboy123):

well, if you need to "determine the equation of free motion", i would have though that you need to derive the differential equation that you mention. as this is in the vertical plane, you might wish to establish why g does not feature, ie why the DE is the same as for spring-mass system in horizontal plane. then, actually solving the DE, using a standard solution, will involve plugging in initial values for displacement and veclocity and so on.

OpenStudy (anonymous):

Ok I know how to solve the DE and get general solution but I don't know how to write the DE given this information

OpenStudy (irishboy123):

this will guide the derivation is that is what you think you need to do. then indeed you take the boilerplate solution x(t) = A cos (wt + phase) and differentiate to get v(t) etc and plug in values you will get spring constant k from the equilibrium position.

OpenStudy (irishboy123):

sorry, the link i meant to provide http://www.ux1.eiu.edu/~cfadd/1150/15Period/Vert.html with helpful drawings

OpenStudy (anonymous):

displacement is 2 inches (8-6?). mass is 2 lbs.

OpenStudy (irishboy123):

if x(t) = Acos(wt-d) v(t) = ?? a(t) = ?? w = root (k/m)

OpenStudy (anonymous):

The link you sent makes sense but don't know about the rest...

OpenStudy (anonymous):

Acos(wt-d) = height?

OpenStudy (anonymous):

or length?

OpenStudy (anonymous):

Do I take Acos(wt-d) and div once for v(t) and twice for a(t)?

OpenStudy (anonymous):

Do i need to plug in values for A,w,t,d and x(t) before i take those derivatives?

OpenStudy (irishboy123):

x (t) = Acos(wt-d) is displacement from mean position as function of time yes, differentiate for v(t) and again for a(t) but you will need omega which is root (k/m)

OpenStudy (anonymous):

w=root (k/m)

OpenStudy (anonymous):

so I need to set up F=-km to solve for k? But what is F?

OpenStudy (irishboy123):

hint "A mass weighing 2 lbs stretches a spring 6 inches. " get k from that??

OpenStudy (anonymous):

F=Ma or F=-km

OpenStudy (anonymous):

so i set up ma=-km

OpenStudy (anonymous):

m=2 a=? k=? 2a=-2k

OpenStudy (irishboy123):

at equilibrium position mg = kx you're doing this in imperial which is not my game, but **if** g in imperial is 32 then 2 * 32 = k * 6

OpenStudy (irishboy123):

yes, from wiki: g = acceleration of gravity (9.81 m/s2, 32.17405 ft/s2) so that works.

OpenStudy (anonymous):

f=-kx? I thought it was -km lol

OpenStudy (anonymous):

I thought i assume g is turned of or something

OpenStudy (anonymous):

is x the initial stretch i.e. 6 in? because that is when it finally stops and is at rest?

OpenStudy (irishboy123):

in the equilibrium position, gravity and spring forces are equal and opposite. so at that position you can say mg = kx and calculate k for the spring "A mass weighing 2 lbs stretches a spring 6 inches. " => 2 * g = k * 6

OpenStudy (anonymous):

k=32/3 w= 4/sqrt3

OpenStudy (anonymous):

ok ty for that i feel i am almost there

OpenStudy (anonymous):

what is d in the Acos(wt-d)? and is t initial time =0?

OpenStudy (anonymous):

so diff Acos(4/sqrt(3)*t(0)-d)

OpenStudy (anonymous):

or do i leave t alone and that is just the variable

OpenStudy (irishboy123):

it is the phase shift. if everything started at t = 0, x = A, then it would be a pure cosine curve. however, that is not the case. you will need \[x = A \cos (\omega t - \delta); \ \dot x = - \omega A \sin (\omega t - \delta)\] at t = 0, which is where the initial values are, you have \[x(0) = A \cos ( - \delta); \ \dot x(0) = - \omega A \sin ( - \delta); \ - \omega \tan ( - \delta) = \frac{\dot x(0)}{x(0)}\]

OpenStudy (irishboy123):

\[\tan (- \delta) = - \frac{1}{\omega} \times \frac{\dot x(0)}{x(0)} = - \sqrt{\frac{3}{16}} \times \frac{\frac{-4}{5}}{8} = \frac{\sqrt{3}}{40}\] not paying great attention to arithmetic so please check the numbers...we should try and agree

OpenStudy (anonymous):

Thank you so much I get it now!